Prior to their phaseout in the 1980s, chemicals containing lead were commonly added to gasoline as anti-knocking agents. A 5.825 g sample of one such additive containing only lead, carbon, and hydrogen was burned in an oxygen rich environment. The products of the combustion were 6.340 g of CO2(g) and 3.244 g of H2O(g). Insert subscripts below to appropriately display the empirical formula of the gasoline additive: CHPb
Answer #1

we know that

moles = mass / molar mass

so

moles of C02 = 6.34 / 44 = 0.1441

now

moles of C = moles of C02 = 0.1441

now

mass of C = 0.1441 x 12 = 1.7292

now

moles of H2O = 3.244 / 18 = 0.180222

moles of H = 2 x moles of H20

moles of H = 2 x 0.18022 = 0.36044

so

mass of H = 0.36044 x 1 = 0.36044

now

mass of lead = mass of sample - ( mass of C + mass of H)

so

mass of lead = 5.825 - ( 1.7292 + 0.36044)

mass of lead = 3.735355

now

moles of lead = 3.735355 / 207.2

moles of lead = 0.0180277

now

consider the ratio of moles

C:H:Pb = 0.1441 : 0.36044 : 0.0180277

C:H:Pb = 8: 20:1

so

the empirical formula is C8H20Pb